# Fact or Guesswork? Evaluating Large Language Model’s Medical Knowledge with Structured One-Hop Judgment

### v1

@@ Abstract
Large language models (LLMs) have been widely adopted in various downstream task domains. However, their ability to directly recall and apply factual medical knowledge remains under-explored. Most existing medical QA benchmarks assess complex reasoning or multi-hop inference, making it difficult to isolate LLMs’ inherent medical knowledge from their reasoning capabilities. Given the high-stakes nature of medical applications, where incorrect information can have critical consequences, it is essential to evaluate how well LLMs encode, retain, and recall fundamental medical facts.
To bridge this gap, we introduce the Medical Knowledge Judgment Dataset (MKJ), a dataset specifically designed to measure LLMs’ one-hop factual medical knowledge. MKJ is constructed from the Unified Medical Language System (UMLS), a large-scale repository of standardized biomedical vocabularies and knowledge graphs. We frame knowledge assessment as a binary judgment task, requiring LLMs to verify the correctness of medical statements extracted from reliable and structured knowledge sources.
Our experiments reveal that LLMs struggle with factual medical knowledge retention, exhibiting significant performance variance across different semantic categories, particularly for rare medical conditions. Furthermore, LLMs show poor calibration, often being overconfident in incorrect answers. To mitigate these issues, we explore retrieval-augmented generation, demonstrating its effectiveness in improving factual accuracy and reducing uncertainty in medical decision-making.
@@
Fact or Guesswork? Evaluating Large Language Model’s Medical Knowledge with Structured One-Hop Judgment
Jiaxi Li1  Yiwei Wang2  Kai Zhang3  Yujun Cai4 Bryan Hooi5  Nanyun Peng6  Kai-Wei Chang6  Jin Lu1 1University of Georgia  2University of California, Merced 3Lehigh University  4The University of Queensland 5National University of Singapore  6University of California, Los Angeles
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 OpenAI (2023) and Llama3 Dubey et al. (2024) have demonstrated remarkable generative capabilities, achieving state-of-the-art performance in various NLP tasks. However, their ability to accurately recall and apply domain-specific knowledge remains a major challenge, particularly in high-stakes fields such as healthcare Li et al. (2022); Bi et al. (2024); Yang et al. (2023); Liu et al. (2023); Yan et al. (2023). LLMs are prone to hallucinations-generating plausible but incorrect content Ji et al. (2023); Bang et al. (2023); Zhang et al. (2023)-which raises concerns about their factual reliability Chen et al. (2023b); Li et al. (2024). These domain-specific knowledge gaps often lead to contradictions, misinformation, and overconfident yet incorrect predictions, making them unsuitable for direct deployment in clinical decision-making.
@@
To evaluate “how knowledgeable are LLM in medicine and healthcare”, existing benchmarks typically frame the task as a question-answering (QA) challenge, often involving clinical questions that require multi-step reasoning, indirect relationships, or external retrieval Pal et al. (2022); Malaviya et al. (2024). However, evaluating LLMs’ core medical knowledge requires a direct and controlled framework-one that can systematically quantify what LLMs “know” without the confounding effects of reasoning or retrieval augmentation. Fundamental knowledge evaluation is not merely an auxiliary task-it is a prerequisite for trustworthy and effective medical foundation models Zhang et al. (2024); Moor et al. (2023). By ensuring that LLMs possess a solid factual foundation, we pave the way for more reliable reasoning, clinical applications, and ultimately, safer deployment in real-world healthcare settings.
@@
To address this gap, we introduce the Medical Knowledge Judgment Dataset (MKJ), designed to systematically evaluate LLMs’ inherent factual medical knowledge through one-hop binary judgment tasks. We construct MKJ through a systematic process of extracting knowledge triplets from Unified Medical Language System (UMLS) and transforming them into carefully templated judgment questions. We focus on one-hop relation and remove triplets with multi-hop nodes or multiple relationships, which ensures single and indisputable answer Wei et al. ; Sun et al. (2024a). Some examples are provided in Figure 1.
@@
The UMLS serves as the ideal foundation for our dataset due to its unparalleled comprehensiveness and reliability in the medical domain. First, UMLS is a rigorously curated and widely trusted biomedical resource, integrating over 3.8 million concepts and 78 million relationships from authoritative medical terminologies. This ensures that the knowledge assessed in our dataset is clinically validated and standardized. Second, its knowledge graph (KG) structure provides an explicit, structured representation of medical knowledge, enabling precise fact-based evaluation while minimizing ambiguity and inconsistencies often found in alternative sources (Abacha et al., 2017; Malaviya et al., 2024). By leveraging UMLS as a gold-standard knowledge base, our dataset ensures high factual reliability, broad medical coverage, and systematic evaluation of LLMs’ medical expertise.
@@
To comprehensively examine LLMs’ capabilities in medical knowledge retention, we focus on four specific research questions:
  • RQ1 To what extent can LLMs accurately perform medical judgments?
  • RQ2 How well are LLMs calibrated in medical and healthcare contexts?
@@
  • RQ3 What are the underlying reasons behind LLMs’ failure to retain certain critical medical knowledge?
  • RQ4 What strategies can enhance the response accuracy of LLMs?
@@
We groups the questions into three progressive categories: Biomedical Entities (foundational concepts), Pathological Conditions (phenomena), and Clinical Practice (applications) with detailed information in Appendix A.1. Through comprehensive analysis on MKJ, we find that LLMs, especially open-source LLMs, still struggle with basic factual medical knowledge retention as illustrated in Figure 1. Our further investigation in later sections reveals critical challenges: poor calibration with frequent overconfidence in incorrect predictions as shown in Figure 4, and notably degraded performance when handling rare medical conditions as in Figure 5. To address these limitations, we implement retrieval-augmented generation, which substantially improves factual accuracy in medical contexts displayed in Table 3.
@@ 2 Related Work
Benchmarks
Most works that study the factuality of LLMs use existing question-answering (QA) benchmarks such as WebQuestions Berant et al. (2013), TriviaQA Joshi et al. (2017), LC-QuAD Trivedi et al. (2017); Dubey et al. (2019), Natural Questions Kwiatkowski et al. (2019), and EntityQuestions Sciavolino et al. (2021). There are also new QA benchmarks assessing LLMs’ long-tail knowledge factuality Mallen et al. (2023); Kumar et al. (2024). For medical QA datasets, Qiu et al. (2024) build MMedBench, a multilingual medical multi-choice QA benchmark with rationales, which involves the aggregation of various medical multi-choice QA datasets from multiple languages Jin et al. (2021); Kasai et al. (2023); Labrak et al. (2022); Vilares and Gómez-Rodríguez (2019); Blinov et al. (2022). There are also some benchmarks involving more challenging medical questions such as the clinical topics datasets of MMLU Hendrycks et al. (2020) and MedMCQA Pal et al. (2022); Malaviya et al. (2024). However, these datasets mostly contain multi-hop or indirect relationship questions, assessing more about skilled reasoning abilities of LLMs. Compared with these benchmarks, MKJ focuses on fundamental evaluation by one-hop direct medical knowledge judgments and contributes to the growing body of work on LLM evaluation by systematically leveraging knowledge graphs.
LLMs in Medicine
To apply LLMs on different tasks in general medical and healthcare contexts, some approaches are adopted such as prompt engineering Shi et al. (2024b); Chen et al. (2024b); Singhal et al. (2025) and fine-tuning LLMs with domain-specific data Singhal et al. (2025); Xu et al. (2023); Xie et al. (2024); Chen et al. (2023c); Shi et al. (2024c); Xiong et al. (2023); Zhao et al. (2024). There are also some work integrating medical knowledge graphs to assist LLMs on tasks such as medical question answering Yang et al. (2024); Yasunaga et al. (2022) and diagnosis prediction Afshar et al. (2024); Gao et al. (2023). While existing efforts have focused on enhancing LLMs’ performances across diverse tasks, a fundamental assessment of their capabilities to internalize and apply medical knowledge remains relatively limited. To address this gap, we introduce the MKJ dataset.
@@ 3.2 Triplets collection
Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ represent the medical knowledge graph derived from UMLS, where $\mathcal{V}$ is the set of vertices (i.e., medical concepts or entities), and $\mathcal{E}$ is the set of predicates (relations) among those concepts.
@@
1. Preliminary Question Set. We begin with a set of multi-hop medical QA benchmarks $\mathcal{Q}_{ori}$, such a LiveQA Abacha et al. (2017) and ExpertQA Malaviya et al. (2024). Each question $Q\in\mathcal{Q}_{ori}$ is a natural-language query. It provides a diverse set of real-world medical problems and serve as a basis for identifying the most relevant and valuable UMLS knowledge components.
@@ 4.1 Experiment setup
We evaluate representative open-source and closed-source LLMs of various sizes and architectures, with an intention on LLMs with smaller sizes. Closed-source models include GPT-3.5-Turbo, GPT-4o-mini, GPT-4o OpenAI (2023), Claude-3-haiku and Claude-3-sonnet cla . Open-source LLMs include Llama-3 series (including Llama-3-8B, Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3BDubey et al. (2024), Ministral-8B Mistral (2024), Qwen2.5 series (including Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3BTeam (2024), and Phi-3 series (including Phi-3-mini-4k, Phi-3-mini-128k, Phi-3.5-miniAbdin et al. (2024). We also include open-source medical LLMs Meditron-7B Chen et al. (2023c) and MeLlama-13B Xie et al. (2024).
More information can be found in Appendix A.2. For the inputs to LLMs, we utilize zero-shot prompting technique to LLMs for their responses. Detailed prompts can be found in Appendix A.3. We employ the most deterministic settings (i.e., temperature=0 or top_k=1) for all models.
@@ 4.2 Evaluation
Each datapoint in the MKJ dataset contains 1 positive judgments and $k$ negative judgments (we set $k$=3). Therefore, we have three categories of evaluation metrics, positive accuracy $\text{Acc}_{\text{pos}}$ (accuracy on the positive judgment questions), negative accuracy $\text{Acc}_{\text{neg}}$ (accuracy on the negative judgment questions), and accuracy Acc (regular accuracy over all samples in $\text{Acc}_{\text{pos}}$ and $\text{Acc}_{\text{neg}}$ combined), which can be formally defined as:
@@ 4.3 RQ1: To what extent can LLMs accurately perform medical judgments?
Table 1: Averaged results of accuracy metrics for LLMs and their corresponding gaps $\text{Acc}_{\text{gap}}$, measured as the difference between positive ($\text{Acc}_{\text{pos}}$) and negative ($\text{Acc}_{\text{neg}}$) accuracies. All numbers are round to two decimal places.
The results of positive accuracies $\text{Acc}_{\text{pos}}$ and negative accuracies $\text{Acc}_{\text{neg}}$ are shown in Table 1. We also calculate the accuracy gap between $\text{Acc}_{\text{pos}}$ and $\text{Acc}_{\text{neg}}$, which is denoted as $\text{Acc}_{\text{gap}}=\text{Acc}_{\text{pos}}-\text{Acc}_{\text{neg}}$. For LLM families with more than one models tested in our experiment, we report the averaged results (round to two decimal places) for convenience and denote these aggregated values with the suffix -Series (e.g., Llama-3-Series).
@@ 4.4 RQ2: How well are LLMs calibrated in medical and healthcare contexts?
Given the bias in accuracies observed in Section 4.3, we aim to assess the calibration of LLMs Guo et al. (2017); Minderer et al. (2021); Xiao et al. (2022) on the MKJ dataset in this section.
Miscalibration is commonly quantified in terms of Expected Calibration Error (ECE) Naeini et al. (2015), which measures the absolute difference between predictive confidence and accuracy. Here we directly present the calibration curve, which is constructed by plotting the pairs of confidence score and accuracy, instead of further calculating the ECE.
@@
The calibration curve is then constructed by plotting the pairs $\{(\hat{p}_{k},\hat{y}_{k})\}_{k=1}^{M}$. A perfectly calibrated model would have these points lie on the diagonal $y=x$.
@@
The LLMs tested in this study illustrate varying degrees of calibration. Across the models, we can observe that general-domain LLMs exhibit significant disparities from the perfect calibration line, while medical LLMs demonstrate better calibration performances.
For general-domain LLMs Ministral-8B, Qwen2.5-3B, and Phi-3.5-mini all reveal erratic status, exhibiting both overconfidence and underconfidence patterns. For Llama-3.2-3B, it is observed that although the model occasionally approaches the ideal calibration line when confidence is low, its uncertainty level is still high when the confidence increases.
For medical LLMs, the calibration of Meditron-7B is far from satisfactory, particularly in the mid-confidence range (0.4–0.8), where it tends to be overconfident. Its accuracy improves at higher confidence scores but still remains inconsistent. As for MeLlama-13B, the curve shows better calibration overall, with predictions closely aligned with the perfect calibration in mid-to-high confidence range (0.5-1.0) although it still tends to be overconfident at low confidence levels (<0.5).
Overall, our analysis reveals pervasive calibration deficiencies across all evaluated models on the MKJ dataset. Poor calibration can result in overconfidence, which may lead to misplaced trust in incorrect predictions Xiong et al. (2024); Chen et al. (2023a), or underconfidence, which could cause clinicians to overlook valuable insights Yang and Thompson (2010); Jiang et al. (2012). This points to calibration as a critical area for LLMs in healthcare and medicine.
@@
Table 3: Performances for RAG of LLMs with sparse and dense retrieval methods. All numbers are in percentage (%). Highest accuracies are in bold, and worst results are marked with underline. The values in parentheses represent the absolute difference between RAG and zero-shot performances.
Our analysis indicates that the performance degradation primarily stems from the presence of rare medical entities within these categories. Through comprehensive cross-referencing with medical databases and literature Chen et al. (2024c); Wang et al. (2024), we identify the prevalence of rare medical conditions as a significant factor contributing to LLM failures. This finding is particularly evident in the most error-prone cases, which include uncommon medical phenomena such as “Swallow syncope” and “Asterixis” within the semantic type Sign or Symptom, as well as rare pharmaceutical compounds like “Prednisolone hexanoate” and “Fluprednisolone” in the semantic type Hormone.
To further quantify the rarity of the medical terms encountered, we leverage the PubMed API [^2] https://www.ncbi.nlm.nih.gov/books/NBK25497/ to analyze their prevalence in the medical literature. Specifically, we calculate the averaged frequencies of the terms in the questions per semantic type as shown in Figure 5, from which we can observe certain correlations between the accuracies and frequencies, where semantic types with low accuracies are more likely to have low frequencies.
This finding indicates a key limitation in their training data coverage, highlighting performance disparities across semantic categories due to under-representation of rare medical concepts.

### v2

# Fact or Guesswork? Evaluating Large Language Models’ Medical Knowledge with Structured One-Hop Judgments

@@ Abstract
Large language models (LLMs) have been widely adopted in various downstream task domains. However, their abilities to directly recall and apply factual medical knowledge remains under-explored. Most existing medical QA benchmarks assess complex reasoning or multi-hop inference, making it difficult to isolate LLMs’ inherent medical knowledge from their reasoning capabilities. Given the high-stakes nature of medical applications, where incorrect information can have critical consequences, it is essential to evaluate the factuality of LLMs to retain medical knowledge.
To address this challenge, we introduce the Medical Knowledge Judgment Dataset (MKJ), a dataset derived from the Unified Medical Language System (UMLS), a comprehensive repository of standardized biomedical vocabularies and knowledge graphs. Through a binary classification framework, MKJ evaluates LLMs’ grasp of fundamental medical facts by having them assess the validity of concise, one-hop statements, enabling direct measurement of their knowledge retention capabilities.
Our experiments reveal that LLMs have difficulty accurately recalling medical facts, with performances varying substantially across semantic types and showing notable weakness in uncommon medical conditions. Furthermore, LLMs show poor calibration, often being overconfident in incorrect answers. To mitigate these issues, we explore retrieval-augmented generation, demonstrating its effectiveness in improving factual accuracy and reducing uncertainty in medical decision-making.
@@
Fact or Guesswork? Evaluating Large Language Models’ Medical Knowledge with Structured One-Hop Judgments
Jiaxi Li1  Yiwei Wang2  Kai Zhang3  Yujun Cai4 Bryan Hooi5  Nanyun Peng6  Kai-Wei Chang6  Jin Lu1 1University of Georgia  2University of California, Merced 3Lehigh University  4The University of Queensland 5National University of Singapore  6University of California, Los Angeles plusnli.github.io/med-knowledge-judgment.github.io
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 OpenAI (2023) and Llama3 Dubey et al. (2024) have demonstrated remarkable generative capabilities, achieving state-of-the-art performance in various NLP tasks. However, their abilities to accurately recall and apply domain-specific knowledge remain a major challenge, particularly in high-stakes fields such as medicine and healthcare Li et al. (2022); Yang et al. (2023); Liu et al. (2023); Yan et al. (2023). LLMs are prone to hallucinations - generating plausible but incorrect content Bang et al. (2023); Zhang et al. (2023); Liu et al. (2025), which raises concerns about their factuality Li et al. (2024); Latif et al. (2024). These domain-specific knowledge gaps often lead to contradictions, misinformation, and overconfident yet incorrect predictions, making them unsuitable for direct deployment in clinical decision-making.
@@
To evaluate “how knowledgeable are LLMs in medicine and healthcare”, existing benchmarks typically frame the task as a question-answering (QA) challenge, often involving clinical questions that require multi-step reasoning, indirect relationships, or external retrieval Pal et al. (2022); Malaviya et al. (2024); Zhou et al. (2025). However, evaluating LLMs’ core medical knowledge requires a direct and controlled framework, one that can systematically quantify what LLMs “know” without the confounding effects of reasoning or retrieval augmentation. Fundamental knowledge evaluation is not merely an auxiliary task. It is a prerequisite for trustworthy and effective medical foundation models Zhang et al. (2024); Moor et al. (2023). By ensuring that LLMs possess a solid factual foundation, we pave the way for more reliable reasoning, clinical applications, and ultimately, safer deployment in real-world healthcare settings.
@@
To bridge this gap, we introduce the Medical Knowledge Judgment Dataset (MKJ), which is designed to systematically evaluate LLMs’ inherent factual medical knowledge through one-hop binary judgment task. We construct MKJ through a systematic process of extracting knowledge triplets from Unified Medical Language System (UMLS) and transforming them into carefully templated judgment questions. We focus on one-hop relation and remove triplets with multi-hop nodes or multiple relationships, which ensures single and indisputable answer Wei et al. (2024); Sun et al. (2024a). Some examples are provided in Figure 1.
@@
The UMLS serves as the ideal foundation for our dataset due to its unparalleled comprehensiveness and reliability in the medical domain. First, UMLS is a rigorously curated and widely trusted biomedical resource, integrating over 3.8 million concepts and 78 million relationships from authoritative medical terminologies, which ensures the knowledge assessed in our dataset is clinically validated and standardized. Second, its knowledge graph (KG) structure provides an explicit, structured representation of medical knowledge, enabling precise fact-based evaluation while minimizing ambiguity and inconsistencies often found in alternative sources (Abacha et al., 2017; Malaviya et al., 2024). By leveraging UMLS as a gold-standard knowledge base, we ensure high factual reliability, broad medical coverage, and systematic evaluation of LLMs’ medical factuality. The generated questions are then grouped into three progressive categories: Biomedical Entities (foundational concepts), Pathological Conditions (phenomena), and Clinical Practice (applications) with detailed information in Appendix A.1.
@@
To comprehensively examine LLMs’ capabilities in medical knowledge retention, we focus on four specific research questions (RQs):
  • RQ1: To what extent can LLMs accurately perform medical judgments?
  • RQ2: How well are LLMs calibrated in medical and healthcare contexts?
@@
  • RQ3: What are the underlying reasons behind LLMs’ failure to retain certain critical medical knowledge?
  • RQ4: What strategies can enhance the response accuracy of LLMs?
@@
Through comprehensive analysis on MKJ, we find that LLMs, especially open-source LLMs, still struggle with basic factual medical knowledge retention as illustrated in Figure 1. Our investigation in later sections reveals critical challenges: poor calibration with frequent overconfidence in incorrect predictions as shown in Figure 4, and notably degraded performance when handling rare medical conditions as shown in Figure 5. To address these limitations, we implement retrieval-augmented generation, which substantially improves factual accuracy as displayed in Table 3.
Our work’s contributions extend beyond the introduction of the MKJ dataset. The systematic evaluation and analysis provide comprehensive insights into LLMs’ handling of medical knowledge, illuminating challenges and opportunities that can inform future medical applications.
@@ 2 Related Work
Factuality Benchmarks.
Evaluating LLM factuality presents a non-trivial challenge and various benchmarks are proposed in the general domain Joshi et al. (2017); Trivedi et al. (2017); Dubey et al. (2019); Lin et al. (2022); Sun et al. (2024b); Kim et al. (2023). Recently, the SimpleQA series of works Wei et al. (2024); Cheng et al. (2025); Gu et al. (2025); He et al. (2024) facilitate factuality evaluation by constraining the scope to short, fact-seeking questions with single answers, making factuality assessment more tractable compared to previous long and open-ended model outputs.
Despite the progress, the factuality evaluation in medicine and healthcare for LLMs nowadays is still a lack to some extent. There are some prior work exploring direct probing of medical knowledge in language models Meng et al. (2022); Zhou et al. (2024). However, they either rely on cloze-style queries that may inadvertently assess reasoning rather than pure knowledge, or employ LLM-based generation methods that could introduce noise and bias. In contrast, our MKJ dataset employs binary classification questions with a systematic construction pipeline based on entity substitution, providing a more direct and unambiguous evaluation of medical knowledge in LLMs.
LLMs in Medicine.
To apply LLMs on different tasks in medical and healthcare contexts, some approaches are adopted such as prompt engineering Shi et al. (2024b); Chen et al. (2024b); Singhal et al. (2025) and fine-tuning LLMs with domain-specific data Xu et al. (2023); Xie et al. (2024); Chen et al. (2023b); Shi et al. (2024c); Xiong et al. (2023); Zhao et al. (2024). There are also some work integrating medical knowledge graphs to assist LLMs on tasks such as medical question answering Yang et al. (2024); Yasunaga et al. (2022) and diagnosis prediction Afshar et al. (2024); Gao et al. (2023). While existing efforts have focused on enhancing LLMs’ performances across various tasks, fundamental assessments of their capabilities to internalize and apply medical knowledge remain relatively limited. Our proposed MKJ dataset bridges this gap by constructing questions covering various factual medical knowledge types, and answers in the easily verifiable short-form format.
@@ 3.2 Triplets collection
Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ represent the medical knowledge graph derived from UMLS, where $\mathcal{V}$ is the set of vertices (i.e., medical concepts or entities), and $\mathcal{E}$ is the set of predicates (relations) among them.
@@
1. Preliminary Question Set. We begin with a set of multi-hop medical QA benchmarks $\mathcal{Q}_{ori}$, such LiveQA Abacha et al. (2017) and ExpertQA Malaviya et al. (2024). Each question $Q\in\mathcal{Q}_{ori}$ is a natural-language query. It provides a diverse set of real-world medical problems and serve as a basis for identifying the most relevant and valuable UMLS knowledge components.
@@ 4.1 Experiment setup
We evaluate representative open-source and closed-source LLMs of various sizes and architectures, with an intention on LLMs with smaller sizes. Closed-source models include GPT-3.5-Turbo, GPT-4o-mini, GPT-4o OpenAI (2023), Claude-3-haiku and Claude-3-sonnet cla . Open-source LLMs involve Llama-3 series (including Llama-3-8B, Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3BDubey et al. (2024), Ministral-8B Mistral (2024), Qwen2.5 series (Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3BTeam (2024), Phi-3 series (including Phi-3-mini-4k, Phi-3-mini-128k, Phi-3.5-miniAbdin et al. (2024), as well as medical LLMs Meditron-7B Chen et al. (2023b) and MeLlama-13B Xie et al. (2024).
More information can be found in Appendix A.2. We utilize zero-shot prompting technique to LLMs for their responses and employ the most deterministic setting (i.e., temperature=0 or top_k=1). Detailed prompts can be found in Appendix A.3.
@@ 4.2 Evaluation
Each datapoint in the MKJ dataset contains 1 positive judgments and $k$ negative judgments (we set kk=3). Therefore, we have three categories of evaluation metrics, positive accuracy $\text{Acc}_{\text{pos}}$ (accuracy on the positive judgment questions), negative accuracy $\text{Acc}_{\text{neg}}$ (accuracy on the negative judgment questions), and accuracy Acc (regular accuracy over all samples in $\text{Acc}_{\text{pos}}$ and $\text{Acc}_{\text{neg}}$ combined), which can be formally defined as:
@@ 4.3 RQ1: To what extent can LLMs accurately perform medical judgments?
Table 1: Averaged results of accuracy metrics for LLMs and their corresponding gaps $\text{Acc}_{\text{gap}}$, measured as the difference between positive ($\text{Acc}_{\text{pos}}$) and negative ($\text{Acc}_{\text{neg}}$) accuracies.
The results of positive accuracies $\text{Acc}_{\text{pos}}$ and negative accuracies $\text{Acc}_{\text{neg}}$ are shown in Table 1. We also calculate the accuracy gap between $\text{Acc}_{\text{pos}}$ and $\text{Acc}_{\text{neg}}$, which is denoted as $\text{Acc}_{\text{gap}}=\text{Acc}_{\text{pos}}-\text{Acc}_{\text{neg}}$. For LLM families with more than one models tested in our experiment, we report the averaged results (round to two decimal places) for convenience and denote these aggregated values with the suffix -Series (e.g., Llama-3-Series). The full results for all models are displayed in Table 5.
@@ 4.4 RQ2: How well are LLMs calibrated in medical and healthcare contexts?
Given the bias in accuracies observed in Section 4.3, we aim to assess the calibration of LLMs Guo et al. (2017); Minderer et al. (2021); Xiao et al. (2022) in this section.
Miscalibration is commonly quantified in terms of Expected Calibration Error (ECE) Naeini et al. (2015), which measures the absolute difference between predictive confidence and accuracy. For better visualization, we present the calibration curve by plotting the pairs of confidence and accuracy.
@@
The calibration curve is then presented by plotting $\{(\hat{p}_{k},\hat{y}_{k})\}_{k=1}^{M}$. A perfectly calibrated model would have these points lie on the diagonal $y=x$.
@@
The LLMs tested in this study illustrate varying degrees of calibration. We can observe that general-domain LLMs exhibit significant disparities from the perfect calibration line, while medical LLMs demonstrate better calibration performances.
For general-domain LLMs, they reveal erratic status, exhibiting both overconfidence and underconfidence patterns. It is especially notable for Phi-3.5-mini, which may be due to its training process as other models in Phi-3-Series also observed similar patterns as shown in Figure  11. For Llama-3.2-3B, although it occasionally approaches the ideal calibration line when confidence is low, its uncertainty level is still high when the confidence increases.
For medical LLMs, the calibration of Meditron-7B is far from satisfactory, particularly in the mid-confidence range (0.4-0.8), where it tends to be overconfident. Its accuracy improves at higher confidence scores but still remains inconsistent. As for MeLlama-13B, the curve shows better calibration overall, with predictions closely aligned with the perfect calibration in mid-to-high confidence range although it still tends to be overconfident at low confidence levels.
Overall, our analysis reveals pervasive calibration deficiencies across all evaluated models on the MKJ dataset. Poor calibration can result in overconfidence, which may lead to misplaced trust in incorrect predictions Xiong et al. (2024); Chen et al. (2023a); Li et al. (2025), or underconfidence, which could cause clinicians to overlook valuable insights Yang and Thompson (2010); Jiang et al. (2012). This points to calibration as a critical area for LLMs in healthcare and medicine.
@@
Table 3: Performances of LLMs for RAG with sparse and dense retrieval methods. All numbers are in percentage (%). Highest accuracies are in bold, and worst results are with underline. The values in parentheses are the absolute differences between RAG and zero-shot, with $\uparrow$ denoting improvement and $\downarrow$ denoting decrement.
We analyze that the observed performance degradation can be primarily attributed to two fundamental challenges in LLM training: long-tail knowledge distribution Kandpal et al. (2023) and co-occurrence bias Kang and Choi (2023). These issues manifest when medical concepts appear infrequently in training data, making it particularly challenging for LLMs to learn and retain accurate information about rare medical entities.
We validate this hypothesis through two complementary analyses. First, through comprehensive review of medical databases and literature Chen et al. (2024c); Wang et al. (2024), we find that the most error-prone cases consistently involve rare medical conditions. This is especially evident in semantic types like Sign or Symptom, where uncommon terms such as "Swallow syncope" and "Asterixis" show high error rates, and Hormone, where rare pharmaceutical compounds like "Prednisolone hexanoate" and "Fluprednisolone" prove particularly challenging for LLMs.
Second, to quantitatively validate the relationship between concept rarity and model performance, we conduct a systematic examination using the PubMed API [^2] https://www.ncbi.nlm.nih.gov/books/NBK25497/. By calculating the average frequency of medical terms per semantic type in the scientific literature, as illustrated in Figure 5, we reveal a clear correlation between term frequency and model accuracy. Semantic types with lower accuracies consistently demonstrate lower frequencies in the medical literature, providing empirical evidence for the impact of the long-tail knowledge.
These findings highlight a critical problem in current LLMs: their performances are significantly impacted by the uneven distribution of medical knowledge in training data, leading to substantial disparities in accuracy across different semantic types, particularly for rare medical concepts.